7.45g KCl is dissolved in 50 mL water. Calculate the concentration of KCl in terms of molarity, molality and normality.

1
Expert's answer
2021-12-13T17:10:10-0500

Mass of KCl = 7.45 g

Molecular wt of KCl = 39 + 35.5 =74.5 g

Moles of KCl = 74.5 / 7.45 ​= 0.1

Molality = moles / mass of solvent (in kg) ​= 0.1 / 50 × 10−3 = 0.5 m

Molarity = moles / volume of solvent (in L) = 0.1 / 50 × 10−3 = 0.5 M

Normality = number of mole equivalents / volume of solvent (in L) = 0.1 / 50 × 10−3 = 0.5 n